OpenStudy (anonymous):

You roll a blue number cube and a green cube. Find P a number greater than 2 on the blue cube and an odd number on the green cube. A. 1/4 B. 1/3 C. 1/12 D. 1/6

OpenStudy (anonymous):

I'm on it now.

OpenStudy (anonymous):

I'm suck on this one though.

OpenStudy (anonymous):

(stuck)

OpenStudy (triciaal):

P(>2) on blue = (3.4.5.6) = 4/6 = 2/3 P(odd) on green = (1, 3, 5) = 3/6 = 1/2 P event A and P event B when events are independent

OpenStudy (anonymous):

To be honest I dont even know where to begin so just tell me what you think the answer is.

OpenStudy (triciaal):

"the probability that both the one and the other occur is the product of their separate probabilities" so 2/3 * 1/2 = 1/3

OpenStudy (anonymous):

ok I'll try if, if you think that

OpenStudy (triciaal):

yes

OpenStudy (anonymous):

good job,,, how'd did you know